# Weighted Envy Freeness With Limited Subsidies

### v1

@@ 4 WEF Solutions for Additive Identical Valuations
Let $A$ be the allocation output by Algorithm 2 under additive identical valuations after $m$ iterations. From Lemma 4.3, it can be deduced that $A$ is WEF-able. Together Proposition 4.5 and Proposition 4.6 imply that, to achieve weighted-envy-freeness under identical additive valuations for the allocation computed by Algorithm 2, the required subsidy per agent $j\in N$ is at most $w_{j}\frac{V}{w_{j}}=V$. In combination with Theorem 3.4, the total required subsidy is at most $(n-1)V$. Note that $W\geq nw_{1}$. Therefore, this bound is better than the one proved in Theorem 3.17: $\left(W-w_{i}\right)V\geq\left(n-1\right)w_{i}V\geq\left(n-1\right)V$.
@@
First, sorting the items according the agents’ valuations can be done in $O(m\log(m))$.
Second, the loop in the algorithm runs for $m$ times. where at each iteration $t\in[m]$, finding the set of agents $I^{t}$ and the agent $i^{t}$ within it takes $O(n)$. Also, allocating the item $o_{t}$ to the agent $i^{t}$ takes $O(1)$.
To sum up, Algorithm 2 runs in $O(mn+m\log(m))$. ∎
The upper bound of $(n-1)V$ is tight even for equal entitlements [15]. Interestingly, when either the valuations or the entitlements are identical, the worst-case upper bound depends on $n$, whereas when both valuations and entitlements are different, the bound depends on $W$.
@@ 5 WEF Solutions for binary additive Valuations
Proposition 5.1.
For every $n$ and every weight vector $\mathbf{w}$, there is an instance with $n$ agents with binary valuations in which the required subsidy in any WEF allocation is at least $\frac{W}{w_{2}}-1$.
@@
There is one item. Agents $1$ and $2$ value the item at $1$ and the others at 00. If agent $i\in\{1,2\}$ gets the item, then the other agent $j$ must get subsidy $\frac{w_{j}}{w_{i}}$. To ensure that other agents do not envy $j$’s subsidy, every other agent $k\not\in\{1,2\}$ must get subsidy $\frac{w_{k}}{w_{i}}$. The total subsidy is $\frac{W}{w_{i}}-1$. The subsidy is minimized by giving the item to agent $2$, since $w_{2}\geq w_{1}$. This gives a lower bound of $\frac{W}{w_{2}}-1$. ∎
@@
We now show how to compute a WEF-able allocation where the subsidy given to each agent $i\in N$ is at most $\frac{w_{i}}{w_{1}}V=\frac{w_{i}}{w_{1}}$.
@@
In the case of binary valuations, Algorithm 1 is inefficient in three ways: (1) the maximum-cost matching does not always prioritize agents with higher entitlements, (2) there may be situations where an agent prefers items already allocated in previous iterations, while the agent holding those items could instead take unallocated ones, and (3) the algorithm works only for agents with integer weights. To illustrate these issues, see the following example:
Example 5.2.
@@
Consider two agents with corresponding weights $w_{1}=1$ and $w_{2}=2$, and five items. The valuation functions are outlined below:
@@
One possible scenario is that Algorithm 1 proceeds as follows:
@@
    At the first iteration, agent 1 receives $o_{1}$ and agent 2 receives $o_{2},o_{3}$.
@@
    At the second iteration, agent 1 receives $o_{4}$, as this matching has a total value of $1$, which is the highest possible value at this stage;
@@
    Similarly,at the third iteration, agent 1 receives $o_{5}$.
In this case, the subsidy required by agent 2 is $\left(\frac{2}{1}-\frac{2}{2}\right)2=2$ which is higher than $W-w_{2}=3-2=1$. An allocation requiring less subsidy would be $A_{1}=\{o_{5}\}$ and $A_{2}=\{o_{1},o_{2},o_{3},o_{4}\}$. Here, the subsidy required by agent 1 is $(\frac{4}{2}-1)2=1$. (an even better allocation would be $A_{1}=\{o_{4},o_{5}\}$ and $A_{2}=\{o_{1},o_{2},o_{3}\}$, which is already WEF without any subsidy).
We address them by modifying the General Yankee Swap (GYS) algorithm introduced by Viswanathan et al. in [26].
GYS starts with an initial allocation, denoted $A^{0}$. In this allocation all agents’ bundles are empty, $A_{i}^{0}=\emptyset$, $\forall i\in N$. For convenience, we add a dummy agent $i_{0}$ and assume that all items are initially assigned to $i_{0}$: $A_{i_{0}}^{0}=M$. The algorithm runs in $T$ iterations . For $t\in\{0,\ldots,T\}$, we denote by $A^{t}$ the allocation at the end of the iteration $t$. Throughout this algorithm, we divide the agents into two sets:
  1. 1.
    $U^{t}$: The agents remaining in the game at the beginning of the iteration $t$.
  2. 2.
    $N\setminus U^{t}$: The agents who were removed from the game in earlier iteration $t^{\prime}<t$. Agents are removed from the game when the algorithm deduces that their utility cannot be improved.
As long as not all the objects have been allocated, at every iteration $t\in[T]$, the algorithm looks for the agents maximizing the gain function (Viswanathan et al. in [26]) among $U^{t}$, i.e., the agents remaining in the game at this iteration.
In our setting we use the gain function: $\left(\frac{w_{i}}{v_{i}(A_{i}^{t})+1},1\right)$, which means that we first select the agents with the minimal potential to increase envy towards them (the maximum $\frac{w_{i}}{v_{i}(A_{i}^{t})+1}$), and then break ties in favor of the agent with the higher index, which corresponds to the agent with the larger $w_{i}$.
The selected agent then faces a choice: either acquire an unallocated item or take an item allocated to another agent. Regardless of the choice, the agent’s utility increases by 1. If the agent decides to take an allocated item from another agent, the affected agent must then decide whether to acquire an unallocated item or take an allocated item from a third agent to maintain their utility, and so on. This results in a path from agent $i$ to the dummy agent $i_{0}$, where items are passed between agents until an unallocated item is reached.
We call this a transfer paths. When an agent is selected, the algorithm attempts to find a transfer path from this agent to $i_{0}$, maintaining utilities for all agents except the initiator, whose utility increases by 1. If no such path exists, the agent will no longer receive any items, and the agent is removed from the game.
We use the polynomial-time method by Viswanathan et al. in [26] to find transfer paths.
Algorithm 3 differs from $GYS$ in the following way: at the beginning of iteration $t$, the algorithm first removes all agents without a transfer path originating from them (Phase 1). Then, it selects an agent based on the gain function to allocate a new item to that agent (Phase 2).
1:Instance $(N,M,\{v_{i}\}_{i\in N},\{w_{i}\}_{i\in N})$ with additive binary valuations.
2:WEF-able allocation $A=(A_{i})_{i\in N}$ with total required subsidy of at most $\frac{W}{w_{1}}-1$.
3:$A_{i_{0}}^{0}\leftarrow M$, and $A_{i}^{0}\leftarrow\emptyset$ for each $i\in N$triangleright All items initially are unassigned
4:$U^{1}\leftarrow N$
5:$t\leftarrow 1$
6:while $U^{t}\neq\emptyset$ do
7:
8:Phase 1 – Removing all agents without a transfer path starting from them
9:
10:     for $i\in U^{t}$ do
11:         Find a transfer path starting at $i$
12:         if  no such path exists then
13:              $U^{t}\leftarrow U^{t}\setminus\{i\}$
14:         end if
15:     end for
16:
17:Phase 2 – Updating the resulting allocation
18:
19:     $I^{t}\leftarrow\arg\max_{i\in U^{t}}\left(\frac{w_{i}}{v_{i}(A_{i}^{t-1})+1}\right)$triangleright Choose the agent who maximizes the gain function
20:     $i^{t}\leftarrow\max_{i\in S}\left(i\right)$triangleright Break ties using index
21:     Find the shortest transfer path starting at $i^{t}$
22$\triangleright$ this path must exist, as $i^{t}$ would have been removed otherwise
23:     Transfer the items along the path and update the allocation $A^{t}$
24:     $U^{t+1}\leftarrow U^{t}$
25:     $t\leftarrow t+1$
26:end while
27:return $A^{t}$
Example 5.3.
For Example 5.2, the algorithm is executed as follows:
  1. 1.
    For $t=1$, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{1}$, $\frac{1}{w_{2}}=\frac{1}{2}$. Consequently, the algorithm searches for the shortest transfer path starting at $i_{2}$ and ending at $i_{\text{0}}$, and finds the path $(i_{2},i_{\text{0}})$. The algorithm transfers the item $o_{1}$ to agent $i_{2}$ from $i_{\text{0}}$’s bundle, resulting in $A_{1}^{1}=\emptyset$ and $A_{2}^{1}=\{o_{1}\}$.
  2. 2.
    For $t=2$, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{1}$ and $\frac{v(A_{2}^{1})+1}{w_{2}}=\frac{2}{2}$. Since those values are equal, the algorithm compares $w_{1}=1$ and $w_{2}=2$. Subsequently, the algorithm searches for the shortest transfer path starting at $i_{1}$ and ending at $i_{\text{0}}$, and finds the path $(i_{2},i_{\text{0}})$. The algorithm transfers the item $o_{2}$ to agent $i_{2}$, yielding $A_{1}^{2}=\emptyset$ and $A_{2}^{2}=\{o_{1},o_{2}\}$.
  3. 3.
@@
    For $t=3$, the algorithm compares $\frac{1}{w_{1}}=1$ and $\frac{v_{2}(A_{2}^{2})+1}{w_{2}}=\frac{3}{2}$. As a result, the algorithm searches for the shortest transfer path starting at $i_{1}$ and ending at $i_{\text{0}}$, and finds the path $(i_{1},i_{\text{0}})$. The algorithm transfers the item $o_{3}$ to agent $i_{1}$, producing $A_{1}^{3}=\{o_{3}\}$ and $A_{2}^{3}=\{o_{1},o_{2}\}$.
@@
    For $t=4$, the algorithm compares $\frac{v_{1}(A_{1}^{3})+1}{w_{1}}=2$ and $\frac{v_{2}(A_{2}^{3})+1}{w_{2}}=\frac{3}{2}$. Thus, the algorithm searches for the shortest transfer path starting at $i_{2}$ and ending at $i_{\text{0}}$, and finds the path $(i_{2},i_{\text{0}})$. The algorithm transfers the item $o_{4}$ to agent $i_{2}$, leading $A_{1}^{4}=\{o_{3}\}$ and $A_{2}^{4}=\{o_{1},o_{2},o_{4}\}$.
@@
    For $t=5$, the algorithm compares $\frac{v_{1}(A_{1}^{4})}{w_{1}}=2$ and $\frac{v_{2}(A_{2}^{2})+1}{w_{2}}=2$. Since those values are equal, the algorithm compares $w_{1}=1$ and $w_{2}=2$. As a consequence, the algorithm searches for the shortest transfer path starting at $i_{2}$ and ending at $i_{\text{0}}$, and finds the path $(i_{2},i_{1},i_{\text{0}})$. The algorithm transfers the item $o_{3}$ to agent $i_{2}$ from $i_{1}$’s bundle and the item $o_{5}$ to agent $i_{1}$ from $i_{\text{0}}$’s bundle. The final allocation is $A_{1}^{5}=\{o_{5}\}$ and $A_{2}^{5}=\{o_{1},o_{2},o_{3},o_{4}\}$.
@@
    The envies are stated below:
    $\begin{bmatrix}&i_{1}&i_{2}\\ i_{1}&0&\frac{4}{2}-\frac{1}{1}=1\\ i_{2}&0-\frac{4}{2}=-2&0\\ \end{bmatrix}$
@@
    In order to mitigate envy, $s_{i_{1}}=1$ and $s_{i_{2}}=0$. Note that the agent with second-smallest weight, $i_{2}$, receives a subsidy of 00.
Definition 5.4.
(Viswanathan et al. [26]) An allocation $A$ is said to be non-redundant if for all $i\in N$, we have $v_{i}(o)=|A_{i}|$.
That is, $v_{j}(A_{i})\leq|A_{i}|=v_{i}(A_{i})$ for every $i,j\in N$. Similarly to Lemma 4.3, we can prove the following Lemma.
Lemma 5.5.
@@
$\begin{split}&cost_{A}(C)=\\ &\sum_{k=1}^{r-1}\left(\frac{v_{i_{k}}(A_{i_{k+1}})}{w_{i_{k+1}}}-\frac{v_{i_{k}}(A_{i_{k}})}{w_{i_{k}}}\right)+\frac{v_{i_{r}}(A_{i_{1}})}{w_{i_{1}}}-\frac{v_{i_{r}}(A_{i_{r}})}{w_{i_{r}}}\leq^{\text{due to non-redundancy}}\\ &\sum_{k=1}^{r-1}\left(\frac{v_{i_{k+1}}(A_{i_{k+1}})}{w_{i_{k+1}}}-\frac{v_{i_{k}}(A_{i_{k}})}{w_{i_{k}}}\right)+\frac{v_{i_{1}}(A_{i_{1}})}{w_{i_{1}}}-\frac{v_{i_{r}}(A_{i_{r}})}{w_{i_{r}}}=\\ &\sum_{k=1}^{r}\left(\frac{v_{i_{k}}(A_{i_{k}})}{w_{i_{k}}}-\frac{v_{i_{k}}(A_{i_{k}})}{w_{i_{k}}}\right)=0\end{split}$ (3)
Hence, by condition (b) of Theorem 3.2, $A$ is WEF-able. ∎
The following is an analogue of Observation 4.2:
Observation 5.6.
Let $A$ be any non-redundant allocation. Let $L=(i,\ldots,j)$ be a path in $G_{A,w}$. Then $cost_{A}(L)\leq\frac{|A_{j}|}{w_{j}}-\frac{|A_{i}|}{w_{i}}$.
Lemma 3.1 in [26] shows that the allocation produced by GYS is non-redundant. The same is true for our variant:
Lemma 5.7.
In any iteration $t\in[T]$ of Algorithm 3, the allocation $A^{t}$ is non-redundant.
@@
We prove by induction that at each iteration $t\in[T]$, $A^{t}$ remains non-redundant.
For the base case, $A^{0}$ is an empty allocation and is therefore non-redundant. Now, assume that at iteration $t-1$, $A^{t-1}$ is non-redundant. Let $A^{t}$ be the allocation resulting from iteration $t$.
If $A^{t-1}=A^{t}$, meaning no agent received a new item, the process is complete. Otherwise, let $i^{t}$ the agent who receives new item. Agent $i^{t}$ obtains an item via the transfer path $L=(i^{t}=i_{1},\ldots,i_{k})$. For each $1\leq j<k$, agent $i_{j}$ receives the item $o_{j}$ from the bundle of $i_{j+1}$, given that $v_{i_{j}}(o_{j})=1$. Agent $i_{k}$ receives a new item $o_{k}$ from the bundle of $i_{0}$, with $v_{i_{k}}(o_{k})=1$.
@@
For agents not on the transfer path $L$, their bundles remain unchanged. Thus, for each agent $i\in N$, it holds that $v_{i}(A_{i}^{t})=|A_{i}^{t}|$, confirming that $A^{t}$ is non-redundant. ∎
Our main task is to prove the following:
Theorem 5.8.
Based on Lemma 5.7 and Lemma 5.5 it is established that for every iteration $t\in[T]$, $A^{t}$ is both non-redundant and WEF-able. Therefore, the remaining task is to establish subsidy bounds.
We focus on two groups: $U^{t}$ and $N\setminus{U^{t}}$. The selection rule simplifies setting limits for the first group (Proposition 5.10). However, understanding the dynamics of the second group presents challenges, as the selection rule is not applicable for them. Hence, the primary objective is to determine the bounds for these subsidies (Proposition 5.14).
Lemma 5.9.
Let $j\in N$ be any agent. After Phase 1 at iteration $t$, if $i\in U^{t}$, then $\frac{|A^{t}_{j}|}{w_{j}}-\frac{|A^{t}_{i}|}{w_{i}}\leq\frac{1}{w_{i}}.$
5.9 is proved similarly to the proof of 4.5, based on the selection rule.
@@
Otherwise, let $t^{\prime}\leq t$ be the latest iteration in which $j$ was selected. By the selection rule, $\frac{v_{j}(A_{j}^{t^{\prime}-1})+1}{w_{j}}\leq\frac{v_{i}(A_{i}^{t^{\prime}-1})+1}{w_{i}}$. Then by non-redundancy, $\frac{|A_{j}^{t^{\prime}}|}{w_{j}}=\frac{v_{j}(A_{j}^{t^{\prime}})}{w_{j}}=\frac{v_{j}(A_{j}^{t^{\prime}-1})+1}{w_{j}}\leq\frac{v_{i}(A_{i}^{t^{\prime}-1})+1}{w_{i}}=\frac{v_{i}(A_{i}^{t^{\prime}})}{w_{i}}+\frac{1}{w_{i}}=\frac{|A_{i}^{t^{\prime}}|}{w_{i}}+\frac{1}{w_{i}}$.
@@
From Lemma 5.9, we can conclude the following:
Proposition 5.10.
After Phase 1 at iteration $t$, if $i\in U^{t}$, then $cost_{A^{t}}(\ell_{i}(A^{t}))\leq\frac{1}{w_{i}}$.

### v4

@@ 4 WEF Solutions for Additive Identical Valuations
Let $A$ be the allocation output by Algorithm 2 under additive identical valuations after $m$ iterations. From Section 4, it can be deduced that $A$ is WEF-able. Together Section 4 and Algorithm 2 imply that, to achieve weighted-envy-freeness under identical additive valuations for the allocation computed by Algorithm 2, the required subsidy per agent $j\in N$ is at most $w_{j}\frac{V}{w_{j}}=V$. In combination with Section 3.1, the total required subsidy is at most $(n-1)V$. Note that $W\geq nw_{1}$. Therefore, this bound is better than the one proved in Algorithm 1: $\left(W-w_{i}\right)V\geq\left(n-1\right)w_{i}V\geq\left(n-1\right)V$.
@@
The loop in the algorithm runs for $m$ times. where at each iteration $t\in[m]$, finding the set of agents $I^{t}$ and the agent $i^{t}$ within it takes $O(n)$. Also, allocating the item $o$ to the agent $i^{t}$ takes $O(1)$.
To sum up, Algorithm 2 runs in $O(mn)$. ∎
Note that $W\geq nw_{1}$. Therefore, this bound is better than the one proved in Algorithm 1 for integer weights: $\left(W-w_{i}\right)V\geq\left(n-1\right)w_{i}V\geq\left(n-1\right)V$.
The upper bound of $(n-1)V$ is tight even for equal entitlements (halpern2019fair). Interestingly, when either the valuations or the entitlements are identical, the worst-case upper bound depends on $n$, whereas when both valuations and entitlements are different, the bound depends on $W$.
@@ 5 WEF Solutions for binary additive Valuations
{proposition}
For every $n\geq 2$ and weight vector $\mathbf{w}$, there is an instance with $n$ agents with binary valuations in which the required subsidy in any WEF allocation is at least $\frac{W}{w_{2}}-1$.
@@
Agents $1$ and $2$ value the item at $1$ and the others at 00. If agent $i\in\{1,2\}$ gets the item, then the other agent $j\neq i\in\{1,2\}$ must get subsidy $\frac{w_{j}}{w_{i}}$. To ensure that other agents do not envy $j$’s subsidy, every other agent $k\not\in\{1,2\}$ must get subsidy $\frac{w_{k}}{w_{i}}$. The total subsidy is $\frac{W}{w_{i}}-1$. The subsidy is minimized by giving the item to agent $2$, since $w_{2}\geq w_{1}$. This gives a lower bound of $\frac{W}{w_{2}}-1$. ∎
@@
Below, we show how to compute a WEF-able allocation where the subsidy given to each agent $i\in N$ is at most $\frac{w_{i}}{w_{1}}V=\frac{w_{i}}{w_{1}}$, and the total subsidy is at most $\frac{W}{w_{1}}-1$.
In the case of binary valuations, Algorithm 1 is inefficient in three ways:
  1. 1.
    The maximum-cost matching does not always prioritize agents with higher entitlements.
  2. 2.
@@
    There may be situations where an agent prefers items already allocated in previous iterations, while the agent holding those items could instead take unallocated ones.
  3. 3.
    The algorithm works only for agents with integer weights.
We address these issues by adapting the General Yankee Swap (GYS) algorithm introduced by Viswanathan et al. in [viswanathan2023general].
GYS starts with an empty allocation for all agents. We add a dummy agent $i_{0}$ and assume that all items are initially assigned to $i_{0}$: $A_{i_{0}}=M$.
Algorithm 3 presents our approach for finding a WEF-able allocation with a bounded subsidy. The algorithm runs in $T$ iterations. We denote by $A^{t}$ the allocation at the end of iteration $t$. Throughout this algorithm, we divide the agents into two sets:
  1. 1.
    $R$: The agents remaining in the game at the beginning of the iteration $t$.
  2. 2.
    $N\setminus R$: The agents who were removed from the game in earlier iteration $t^{\prime}<t$. Agents are removed from the game when the algorithm deduces that their utility cannot be improved.
As long as not all the objects have been allocated, at every iteration $t\in[T]$, the algorithm looks for the agents maximizing the gain function (viswanathan2023general) among $R$, i.e., the agents remaining in the game at this iteration.
We use the gain function: $\frac{w_{i}}{v_{i}(A_{i}^{t-1})+1}$, which selects agents with the minimal potential for increasing envy. If multiple agents have the same value, we select one arbitrarily.
The selected agent then chooses either to acquire an unallocated item or take an item from another agent. In either case, their utility increases by 1. If the agent takes an item from another, the affected agent must decide whether to take an unallocated item or another allocated item to preserve their utility, and so on. This process creates a transfer path from agent $i$ to the dummy agent $i_{0}$ , where items are passed until an unallocated item is reached.
Formally, we represent this as a directed graph, where nodes are agents, and an edge $(i,j)$ if and only if there exists an item in $j$’s bundle that $i$ values positively. A transfer path is any directed path in that graph, that ends at the dummy agent $i_{0}$.
When an agent is selected, the algorithm attempts to find a transfer path from that agent, preserving utilities for all agents except the initiator, whose utility increases by 1. If no path is found, the agent is removed from the game. We use the polynomial-time method by viswanathan2023general to find transfer paths.
Algorithm 3 differs from $GYS$ in the following way: at the beginning of iteration $t$, the algorithm first removes all agents without a transfer path originating from them (line 7). Then, it selects an agent based on the gain function to allocate a new item to that agent. For convenience, we denote by $R(t)$ the agents who have a transfer path originating from them at the beginning of iteration $t$ (line 7).
1:Instance $(N,M,v,\mathbf{w})$ with additive binary valuations.
2:WEF-able allocation $A$ with total required subsidy of at most $\frac{W}{w_{1}}-1$.
3:$A_{i_{0}}\leftarrow M$, and $A_{i}^{0}\leftarrow\emptyset$ for each $i\in N$triangleright All items initially are unassigned
4:$t\leftarrow 1$
5:$R\leftarrow N$
6:while $R\neq\emptyset$ do
7:     Remove from $R$ all agents who do not have a transfer path starting from them
8:     $i^{t}\leftarrow\arg\max_{i\in R}\left(\frac{w_{i}}{v_{i}(A_{i}^{t-1})+1}\right)$triangleright Choose the agent who maximizes the gain function
9:     Find a transfer path starting at $i^{t}$triangleright For example, one can use the BFS algorithm to find a shortest path from $i^{t}$ to $i_{0}$.
10:     Transfer the items along the path and update the allocation $A^{t}$
11:     $t\leftarrow t+1$
12:end while
13:return $A^{t}$
The following example demonstrates Algorithm 3:
[…]
@@
Consider two agents with weights $w_{1}=1$ and $w_{2}=2$, and five items. The valuation functions are:
@@
The algorithm is executed as follows:
@@
    For $t=1$, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{1}$, $\frac{1}{w_{2}}=\frac{1}{2}$. Consequently, the algorithm searches for a transfer path starting at $i_{2}$ and ending at $i_{\text{0}}$, and finds the path $(i_{2},i_{\text{0}})$. The algorithm transfers the item $o_{1}$ to agent $i_{2}$ from $i_{\text{0}}$’s bundle, resulting in $A_{1}^{1}=\emptyset$ and $A_{2}^{1}=\{o_{1}\}$.
@@
    For $t=2$, the algorithm compares $\frac{1}{w_{1}}=\frac{1}{1}$ and $\frac{v_{2}(A_{2}^{1})+1}{w_{2}}=\frac{2}{2}$. Since those values are equal, the algorithm arbitrarily selects agent $i_{2}$ and searches for a transfer path starting at $i_{2}$ and ending at $i_{\text{0}}$, and finds the path $(i_{2},i_{\text{0}})$. The algorithm transfers the item $o_{2}$ to agent $i_{2}$, yielding $A_{1}^{2}=\emptyset$ and $A_{2}^{2}=\{o_{1},o_{2}\}$.
@@
    For $t=3$, the algorithm compares $\frac{1}{w_{1}}=1$ and $\frac{v_{2}(A_{2}^{2})+1}{w_{2}}=\frac{3}{2}$. As a result, the algorithm searches for a transfer path starting at $i_{1}$ and ending at $i_{\text{0}}$, and finds the path $(i_{1},i_{\text{0}})$. The algorithm transfers the item $o_{3}$ to agent $i_{1}$, producing $A_{1}^{3}=\{o_{3}\}$ and $A_{2}^{3}=\{o_{1},o_{2}\}$.
@@
    For $t=4$, the algorithm compares $\frac{v_{1}(A_{1}^{3})+1}{w_{1}}=2$ and $\frac{v_{2}(A_{2}^{3})+1}{w_{2}}=\frac{3}{2}$. Thus, the algorithm searches for a transfer path starting at $i_{2}$ and ending at $i_{\text{0}}$, and finds the path $(i_{2},i_{\text{0}})$. The algorithm transfers the item $o_{4}$ to agent $i_{2}$, leading $A_{1}^{4}=\{o_{3}\}$ and $A_{2}^{4}=\{o_{1},o_{2},o_{4}\}$.
@@
    For $t=5$, the algorithm compares $\frac{v_{1}(A_{1}^{4})+1}{w_{1}}=2$ and $\frac{v_{2}(A_{2}^{4})+1}{w_{2}}=\frac{4}{2}=2$. Since those values are equal, the algorithm arbitrarily selects agent $i_{2}$ and searches for a transfer path starting at $i_{2}$ and ending at $i_{\text{0}}$, and finds the path $(i_{2},i_{1},i_{\text{0}})$. The algorithm transfers the item $o_{3}$ to agent $i_{2}$ from $i_{1}$’s bundle and the item $o_{5}$ to agent $i_{1}$ from $i_{\text{0}}$’s bundle, leading $A_{1}^{5}=\{o_{5}\}$ and $A_{2}^{5}=\{o_{1},o_{2},o_{3},o_{4}\}$.
@@
    Agent 1 envies agent 2 by $\frac{4}{2}-\frac{1}{1}=1$, while agent 2 envies agent 1 by $00-\frac{4}{2}<$ 0.
@@
    In order to mitigate envy, $s_{1}=1$ and $s_{2}=0$.
[…]
(viswanathan2023general) An allocation $A$ is said to be non-redundant if for all $i\in N$, we have $v_{i}(A_{i})=|A_{i}|$.
That is, $v_{j}(A_{i})\leq|A_{i}|=v_{i}(A_{i})$ for every $i,j\in N$. Similarly to Section 4, we can prove the following:
{lemma}
@@
$\begin{split}&cost_{A}(C)=\\ &\sum_{k=1}^{r-1}\left(\frac{v_{i_{k}}(A_{i_{k+1}})}{w_{i_{k+1}}}-\frac{v_{i_{k}}(A_{i_{k}})}{w_{i_{k}}}\right)+\frac{v_{i_{r}}(A_{i_{1}})}{w_{i_{1}}}-\frac{v_{i_{r}}(A_{i_{r}})}{w_{i_{r}}}\leq\\ &\sum_{k=1}^{r-1}\left(\frac{v_{i_{k+1}}(A_{i_{k+1}})}{w_{i_{k+1}}}-\frac{v_{i_{k}}(A_{i_{k}})}{w_{i_{k}}}\right)+\frac{v_{i_{1}}(A_{i_{1}})}{w_{i_{1}}}-\frac{v_{i_{r}}(A_{i_{r}})}{w_{i_{r}}}=\\ &\sum_{k=1}^{r}\left(\frac{v_{i_{k}}(A_{i_{k}})}{w_{i_{k}}}-\frac{v_{i_{k}}(A_{i_{k}})}{w_{i_{k}}}\right)=0,\end{split}$ (3)
where the inequality holds due to non-redundancy. Hence, by condition (b) of Section 3.1, $A$ is WEF-able. ∎
Lemma 3.1 in [viswanathan2023general] shows that the allocation produced by GYS is non-redundant. The same is true for our variant: {lemma} At the end of any iteration $t$ of Algorithm 3, the allocation $A^{t}$ is non-redundant.
@@
We prove by induction that at the end of each iteration $t$, $A^{t}$ remains non-redundant.
For the base case, $A^{0}$ is an empty allocation and is therefore non-redundant. Now, assume that at the end of iteration $t-1$, $A^{t-1}$ is non-redundant.
If $A^{t}=A^{t-1}$, meaning no agent received a new item, the process is complete. Otherwise, let $i^{t}$ be the agent who receives new item. Agent $i^{t}$ obtains an item via the transfer path $P=(i^{t}=i_{1},\ldots,i_{k})$. For each $1\leq j<k$, agent $i_{j}$ receives the item $o_{j}$ from the bundle of $i_{j+1}$, given that $v_{i_{j}}(o_{j})=1$. Agent $i_{k}$ receives a new item $o_{k}$ from the bundle of $i_{0}$, with $v_{i_{k}}(o_{k})=1$.
@@
For agents not on the transfer path $P$, their bundles remain unchanged. Thus, for each agent $i\in N$, it holds that $v_{i}(A_{i}^{t})=v_{i}(A_{i}^{t-1})=|A_{i}^{t-1}|=|A_{i}^{t}|$, confirming that $A^{t}$ is non-redundant. ∎
Based on Algorithm 3 it is established that at the end of every iteration $t\in[T]$, $A^{t}$ is WEF-able. The remaining task is to establish subsidy bounds.
We focus on two groups: $R$ and $N\setminus R$.
The selection rule simplifies limit-setting for $R$ and ensures a subsidy bound of 1 (Section 5). However, understanding the dynamics of the second group, $N\setminus R$, presents challenges, as the selection rule is not applicable for them. For an agent $i\in N\setminus R$, we prove a subsidy bound of $w_{i}\cdot\frac{1}{w_{j}}$, for some $j\in R$. In particular, the bound is at most $\frac{w_{i}}{w_{1}}$.
[…]
Let $A$ be any non-redundant allocation. Let $P=(i,\ldots,j)$ be a path in $G_{A,w}$. Then $cost_{A}(P)\leq\frac{|A_{j}|}{w_{j}}-\frac{|A_{i}|}{w_{i}}$.
{lemma}
Let $j\in N$ be any agent, if $i\in R(t)$, then $\frac{|A^{t}_{j}|}{w_{j}}-\frac{|A^{t}_{i}|}{w_{i}}\leq\frac{1}{w_{i}}.$
@@
Otherwise, let $t^{\prime}\leq t$ be the latest iteration in which $j$ was selected. As agents can not be added to $R$ and by the selection rule, $\frac{v_{j}(A_{j}^{t^{\prime}-1})+1}{w_{j}}\leq\frac{v_{i}(A_{i}^{t^{\prime}-1})+1}{w_{i}}$. Then by non-redundancy,
$\frac{|A_{j}^{t^{\prime}}|}{w_{j}}=\frac{v_{j}(A_{j}^{t^{\prime}})}{w_{j}}=\frac{v_{j}(A_{j}^{t^{\prime}-1})+1}{w_{j}}\leq\frac{v_{i}(A_{i}^{t^{\prime}-1})+1}{w_{i}}=\frac{v_{i}(A_{i}^{t^{\prime}})}{w_{i}}+\frac{1}{w_{i}}=\frac{|A_{i}^{t^{\prime}}|}{w_{i}}+\frac{1}{w_{i}}.$
@@
From Section 5, we can conclude the following: {proposition} If $i\in R(t)$, then $\ell_{i}(A^{t})\leq\frac{1}{w_{i}}$.